OpenStudy (anonymous):

If y varies inversely as x, and y = 3 when x = 6, find the value of x when y = 18.

OpenStudy (lgbasallote):

y varies inversely as x so \[y = \frac{k}{x}\] now...y = 3 when x = 6 so \[3 = \frac{k}{6}\] \[k = 18\] we're going to use this for the second situation what is x when y = 18 \[18 = \frac{k}{x}\] \[18 = \frac{18}{x}\]

OpenStudy (lgbasallote):

what do you think is x?

OpenStudy (anonymous):

1?

OpenStudy (lgbasallote):

you got it!

OpenStudy (anonymous):

thankss!!:)